Kevin will go to college in three years. His parents are opening an educational fund for him, depositing $4000 at the start of each quarter. If the fund assesses a 9 ¼ % interest compounded quarterly, how much will Kevin have in his account when he starts college?
Answer #1

Quarterly interest rate, r = 9.25%/4 = 0.023125

Number of payments = 3 * 4 = 12

PMT FV = =*[(1 + r) - 1

4.000 FV = 0.023125 * ((1+0.023125) 12 - 11

FV = 172,972.972972973 * 0.3156620904

FV = $54, 601.0102313514

Kevin will have $54,601.0102313514 in his account when he starts college.

Can you please upvote? Thank You :-)